# Kinetic formulation of compartmental epidemic models

### v1

@@ 3.2.2 Order 1 Equation
$Q_{S}^{(P)}[\mathbf{f}_{0}]$ $=\int_{V}\Big{(}-\eta(t,x)v\cdot(\nabla_{x}f_{I,0})f_{S,0}^{\prime}+\xi(t,x)v^{\prime}\cdot(\nabla_{x}f_{I,0}^{\prime})f_{S,0}^{\prime}$
@@
where $\eta=\eta(t,x)$, $\xi=\xi(t,x)$, and $F^{\prime}=F(v^{\prime})=F(|v^{\prime}|)$. Note that one particular realization of the leading order equilibrium distribution $F$ is $F(v)=\frac{1}{|V|}$. This will be the case whenever $V$ is a hypersphere (a circle, in the two-dimensional case). In this case, the last equation simplifies to $Q_{S}^{(P)}[\mathbf{f}_{0}]=-(\eta+\xi)vF\cdot(\nabla_{x}\rho_{I})\rho_{S}$.
@@ 3.2.3 Order 2 equations
$\partial_{t}f_{j,0}+v\cdot\nabla_{x}f_{j,1}=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\ .$
Integrating on $V$ and using Eqs. (4) using the conservative property of $Q^{(P)}_{j}$, we find
@@ 3.3 The SIRS epidemic model
However, when working with the SIRS model, we are interested in the total number of type S, I, and R individuals (possibly normalized to the total population), and not its average space-density. Therefore, it is natural to define
@@
$(S,I,R)$ is a solution of the SIRS model for constant-in-space $\rho_{j}$ with parameters
@@
We finish this (small) subsection, noting that it is not difficult to generalize all the above processes for more general class transition kernels. In particular, it is possible to choose class transition kernels
$Q_{j}^{(T)}[\mathbf{f}]=f_{j}\left(\Psi_{j}[\mathbf{f}]-\sum_{j}f_{j}\Psi_{j}[\mathbf{f}]\right)\ ,$
where $\Psi_{j}[\mathbf{f}]$ is the fitness (reproductive viability) function for a population at state $\mathbf{f}$ (i.e, the presence of the type $j$ individuals is given by $f_{j}$. In this case, $\dot{f}_{j}=Q_{j}^{(T)}$ is the well-known replicator equation [HS98].
From the discussion in this section, the kinetic formulation is one possible generalization of the replicator equation, including not only space-inhomogeneity but also other variables (in particular, in this work, velocity).
@@ 4 Existence and uniqueness of solutions of the kinetic model
Let $\alpha,\beta,\gamma\in L^{\infty}((0,T);L^{\infty}(\Omega\times V))$, and consider $f^{0}_{j}\in L^{p}(\Omega\times V)$, for $p\geq 2$. Assume $Q^{(P)}\equiv 0$. Then, there is a unique solution $(f_{S},f_{I},f_{R})$ of the kinetic system (3), with $\varepsilon>0$, with $f_{j}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$, with initial condition given by $(f_{S}^{0},f_{I}^{0},f_{R}^{0})$.
@@
We assume without loss of generality that $\varepsilon=1$ and start with the existence part. Whenever we refer to a constant $C$ (with or without index), we consider that “there exists a positive number $C$ such that”. Its precise value is immaterial to the proof of the theorem, and the same parameter $C$ may refer to different constants along the proof.
Using the boundary condition $\hat{n}\cdot\nabla f_{j}|_{\partial\Omega}=0$, we conclude that
$0=\int_{\Omega}\nabla_{x}f_{j}^{p}\mathrm{d}x=p\int_{\Omega}f_{j}^{p-1}\nabla_{x}f_{j}\mathrm{d}x\ .$
@@
$\frac{1}{p}\frac{\mathrm{d}\ }{\mathrm{d}t}\int_{\Omega\times V}f_{j}^{p}\mathrm{d}v\mathrm{d}x=\int_{\Omega\times V}f_{j}^{p-1}\left[\partial_{t}f_{j}+v\cdot\nabla_{x}f_{j}\right]\mathrm{d}v\mathrm{d}x=\int_{\Omega\times V}f_{j}^{p-1}\mathfrak{Q}_{j}[\mathbf{f}]\mathrm{d}v\mathrm{d}x$ (12)
@@
$\int_{V}f_{j}\mathrm{d}v\leq|V|^{1-\frac{1}{p}}\left(\int_{V}f_{j}^{p}\mathrm{d}v\right)^{1/p}\ .$
@@
$\int_{V}f_{j}^{p-1}\mathrm{d}v\leq|V|^{\frac{1}{p}}\left(\int_{V}f_{j}^{p}\mathrm{d}v\right)^{1-\frac{1}{p}}\ .$
Multiplying the last two equations, we conclude
@@
$=\int_{\Omega}\lambda_{j}\left[\frac{1}{|V|}\int_{V}f_{j}\mathrm{d}v\int_{V}f_{j}^{p-1}\mathrm{d}v-\int_{V}f_{j}^{p}\mathrm{d}v\right]\mathrm{d}x\leq 0\ .$ (13)
We proceed to prove the existence part of $f_{I}$. Note that
$\int_{\Omega\times V}Q^{(T)}_{I}f_{I}^{p-1}\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\left[\frac{\beta f_{S}f_{I}^{p}}{f_{S}+f_{S}+f_{R}}-\gamma f_{I}^{p}\right]\mathrm{d}v\mathrm{d}x$
$\leq\|\beta\|_{L^{\infty}(\Omega\times V)}\|f_{I}\|_{L^{p}(\Omega\times V)}^{p}\ .$
Therefore
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{I}\|_{L^{p}(\Omega\times V)}\leq C\|f_{I}\|_{L^{p}(\Omega\times V)}\ .$
Using Gronwall’s inequality, we conclude that $f_{I}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$.
The next step is to show that
$\int_{\Omega\times V}Q^{(T)}_{R}f_{R}^{p-1}\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\left[\gamma f_{I}f_{R}^{p-1}-\alpha f_{R}^{p}\right]\mathrm{d}v\mathrm{d}x$
$\leq\|\gamma\|_{L^{\infty}(\Omega\times V)}\|f_{I}\|_{L^{p}(\Omega\times V)}\|f_{R}\|_{L^{p}(\Omega\times V)}^{p-1}\ .$
Using the estimate given by Eq. (12) and the existence result for $f_{I}$, we conclude
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{R}\|_{L^{p}(\Omega\times V)}\leq C\ ,$
and, therefore $f_{R}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$.
@@
$\int_{\Omega\times V}Q^{(T)}_{S}[\mathbf{f}]f_{S}^{p-1}\mathrm{d}v\mathrm{d}x\leq\|\alpha\|_{L^{\infty}(\Omega\times V)}\|f_{S}\|_{L^{p}(\Omega\times V)}^{p-1}\ ,$
and, therefore, $f_{S}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$.
Now, we proceed to prove the uniqueness of the solution.
Let us define the norm
$\|\mathbf{f}-\mathbf{g}\|^{2}:=(f_{S}-g_{S})^{2}+(f_{I}-g_{I})^{2}+(f_{R}-g_{R})^{2}\ .$ (14)
From the linearity of $Q^{(R)}_{j}$ it is not difficult to prove that
$Q^{(R)}_{j}[\mathbf{f}]-Q^{(R)}_{j}[\mathbf{g}]=Q^{(R)}_{j}[\mathbf{f}-\mathbf{g}]\ .$
@@
$\int_{\Omega\times V}\left(Q^{(R)}_{j}[\mathbf{f}]-Q^{(R)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)\mathrm{d}v\mathrm{d}x=\int_{\Omega\times V}Q^{(R)}_{j}[\mathbf{f}-\mathbf{g}](f_{j}-g_{j})\leq 0\ .$ (15)
In the sequel, we will use the inequality $ab\leq\frac{1}{2}(a^{2}+b^{2})$.
@@
$\left(Q^{(T)}_{S}[\mathbf{f}]-Q^{(T)}_{S}[\mathbf{g}]\right)\left(f_{S}-g_{S}\right)+\left(Q^{(T)}_{I}[\mathbf{f}]-Q^{(T)}_{I}[\mathbf{g}]\right)\left(f_{I}-g_{I}\right)$
$\qquad+\left(Q^{(T)}_{R}[\mathbf{f}]-Q^{(T)}_{R}[\mathbf{g}]\right)\left(f_{R}-g_{R}\right)$
@@
We conclude that
$\sum_{j=S,I,R}\left(Q^{(T)}_{j}[\mathbf{f}]-Q^{(T)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)$ $\leq C_{1}(f_{S}-g_{S})^{2}+C_{2}(f_{I}-g_{I})^{2}+C_{3}(f_{R}-g_{R})^{2}$
$\leq C\|\mathbf{f}-\mathbf{g}\|^{2}\ .$
Considering the last equation and Eq, (15), we prove that
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|\mathbf{f}-\mathbf{g}\|^{2}\leq 0\ ,$
and the uniqueness follows. ∎
@@
As discussed at the beginning of this section, we did not consider the preferred kernel in the proof of Thm. 2. As will be discussed in the conclusions, we believe that the existence and uniqueness theorem still holds with parameters $\eta,\xi>0$. If Thm. 2 could be extended such that $f_{I}\in L^{\infty}((0,T);H^{1}(\Omega\times V))$, then a bound would follow immediately.
@@
5 Examples
In this section, we consider three different examples. The first one, in Subsec. 5.1, is of a pedagogical nature and deals with a situation with constant parameters.
However, the reason to consider kinetic formulations of epidemic models is to introduce the velocity as a parameter, and this has interest only if we want to study the interaction between the pathogen dynamics and the mobility of individuals.

### v2

@@ 3.2.2 Order 1 Equation
$\mathcal{Q}_{S}^{(P)}[\mathbf{f}_{0}]$ $=\int_{V}\Big(-\eta(t,x)v\cdot(\nabla_{x}f_{I,0})f_{S,0}^{\prime}+\xi(t,x)v^{\prime}\cdot(\nabla_{x}f_{I,0}^{\prime})f_{S,0}^{\prime}$
@@
where $\eta=\eta(t,x)$, $\xi=\xi(t,x)$, and $F^{\prime}=F(v^{\prime})=F(|v^{\prime}|)$. Note that one particular realisation of the leading order equilibrium distribution $F$ is $F(v)=\frac{1}{|V|}$. This will be the case whenever $V$ is a hypersphere (a circle, in the two-dimensional case). In this case, the last equation simplifies to $\mathcal{Q}_{S}^{(P)}[\mathbf{f}_{0}]=-(\eta+\xi)vF\cdot(\nabla_{x}\rho_{I})\rho_{S}$.
@@ 3.2.3 Order 2 equations
$\partial_{t}f_{j,0}+v\cdot\nabla_{x}f_{j,1}=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]+\overline{\mathcal{Q}}^{(P)}_{j}[\mathbf{f}_{0},\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\ .$
Integrating on $V$ and using Eqs. (5) using the conservative property of $\overline{\mathcal{Q}}^{(P)}_{j}$, we find
@@ 3.3 The SIRS epidemic model
However, when working with the SIRS model, we are interested in the total number of type S, I, and R individuals (possibly normalised to the total population), and not its average space-density. Therefore, it is natural to define
@@
We find that $(S,I,R)$ is a solution of the SIRS model for constant-in-space $\rho_{j}$ with parameters
@@
We finish this subsection, noting that it is not difficult to generalise all the above processes for more general class transition kernels. In particular, it is possible to choose class transition kernels
$\mathcal{Q}_{j}^{(T)}[\mathbf{f}]=f_{j}\left(\Psi_{j}[\mathbf{f}]-\sum_{j}f_{j}\Psi_{j}[\mathbf{f}]\right)\ ,$
where $\Psi_{j}[\mathbf{f}]$ is the fitness (reproductive viability) function for a population at state $\mathbf{f}$ (i.e, the presence of the type $j$ individuals is given by $f_{j}$. In this case, $\dot{f}_{j}=\mathcal{Q}_{j}^{(T)}[\mathbf{f}]$ is the well-known replicator equation [HS98].
From the discussion in this section, the kinetic formulation is one possible generalisation of the replicator equation, including not only space-inhomogeneity but also other variables (in particular, in this work, velocity).
@@ 4 Existence and uniqueness of solutions of the kinetic model
Let $\alpha,\beta,\gamma\leq C$, and consider $f^{\mathrm{ini}}_{j}\in L^{p}(\Omega\times V)$, for $p\geq 2$. Assume $\mathcal{Q}^{(P)}_{j}\equiv 0$, $j=S,I,R$. Then, for $\varepsilon>0$, there exists a global solution $(f_{S},f_{I},f_{R})\in\left[L^{\infty}(\mathbb{R}_{+};L^{p}(\Omega\times V))\right]^{3}$ of the kinetic system (3), with initial condition given by $(f_{S}^{\mathrm{ini}},f_{I}^{\mathrm{ini}},f_{R}^{\mathrm{ini}})$.
@@
We assume without loss of generality that $\varepsilon=1$. Whenever we refer to a constant $C$ (with or without index), we consider that “there exists a positive number $C$ such that”. Its precise value is immaterial to the proof of the theorem, and the same parameter $C$ may refer to different constants along the proof.
@@
$\frac{1}{p}\frac{\mathrm{d}\ }{\mathrm{d}t}\int_{\Omega\times V}f_{j}^{p}\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}f_{j}^{p-1}\left[\partial_{t}f_{j}+v\cdot\nabla_{x}f_{j}\right]\mathrm{d}v\mathrm{d}x$ (13)
$=\int_{\Omega\times V}f_{j}^{p-1}\mathfrak{Q}_{j}[\mathbf{f}]\mathrm{d}v\mathrm{d}x$
@@
$\int_{V}f_{j}\mathrm{d}v\leq|V|^{1-\frac{1}{p}}\left(\int_{V}f_{j}^{p}\mathrm{d}v\right)^{1/p}\ .$ (14)
@@
$\int_{V}f_{j}^{p-1}\mathrm{d}v\leq|V|^{\frac{1}{p}}\left(\int_{V}f_{j}^{p}\mathrm{d}v\right)^{1-\frac{1}{p}}\ .$ (15)
We study each term in the right-hand-side of Eq. (13), using that $\mathfrak{Q}_{j}=\mathcal{Q}^{(R)}_{j}+\mathcal{Q}^{(T)}_{j}$. For $\mathcal{Q}^{(R)}_{j}$, we find
@@
$=\int_{\Omega}\lambda_{j}\left[\frac{1}{|V|}\int_{V}f_{j}\mathrm{d}v\int_{V}f_{j}^{p-1}\mathrm{d}v-\int_{V}f_{j}^{p}\mathrm{d}v\right]\mathrm{d}x$
$\leq 0\ ,$ (16)
where the last inequality follows from the multiplication of Inequalities (14) and (15). For $\mathcal{Q}^{(T)}_{I}$, we find
$\int_{\Omega\times V}f_{I}^{p-1}\mathcal{Q}^{(T)}_{I}[\mathbf{f}]\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\left[\frac{\beta f_{S}f_{I}^{p}}{f_{S}+f_{I}+f_{R}}-\gamma f_{I}^{p}\right]\mathrm{d}v\mathrm{d}x$
$\leq C\|f_{I}\|_{L^{p}(\Omega\times V)}^{p}\ ,$ (17)
where we used that $\nicefrac{{f_{S}}}{{(f_{S}+f_{I}+f_{R})}}\leq 1$. Therefore, from Eqs. (13), (16), with $j=I$, and Eq. (17), we find
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{I}\|_{L^{p}(\Omega\times V)}\leq C\|f_{I}\|_{L^{p}(\Omega\times V)}\ \quad\forall t.$
Using Gronwall’s inequality [Gro19], we conclude that $f_{I}\in L^{\infty}(\mathbb{R}_{+};L^{p}(\Omega\times V))$.
Using Holder’s inequality, with $g=f_{I}$, $h=f_{R}^{p-1}$, $q_{1}=p$, $q_{2}=\nicefrac{{p}}{{(p-1)}}$, we show that
$\int_{\Omega\times V}f_{R}^{p-1}\mathcal{Q}^{(T)}_{R}[\mathbf{f}]\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\left[\gamma f_{I}f_{R}^{p-1}-\alpha f_{R}^{p}\right]\mathrm{d}v\mathrm{d}x$
$\leq\int_{\Omega\times V}\gamma f_{I}f_{R}^{p-1}\mathrm{d}v\,\mathrm{d}x$
$\leq C\|f_{I}\|_{L^{p}(\Omega\times V)}\|f_{R}\|_{L^{p}(\Omega\times V)}^{p-1}\ .$ (18)
Using Eqs. (13), (16),  (18), and the existence result for $f_{I}$, we conclude
$\|f_{R}\|^{p-1}_{L^{p}(\Omega\times V)}\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{R}\|_{L^{p}(\Omega\times V)}=\frac{1}{p}\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{R}\|^{p}_{L^{p}(\Omega\times V)}\leq C\|f_{R}\|^{p-1}_{L^{p}(\Omega\times V)}\ ,$
and, therefore,
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{R}\|_{L^{p}(\Omega\times V)}\leq C\ ,\quad\forall t.$
We conclude that $f_{R}\in L^{\infty}(\mathbb{R}_{+};L^{p}(\Omega\times V))$.
@@
$\int_{\Omega\times V}f_{S}^{p-1}\mathcal{Q}^{(T)}_{S}[\mathbf{f}]\mathrm{d}v\mathrm{d}x\leq C\|f_{S}\|_{L^{p}(\Omega\times V)}^{p-1}\ ,$
and, therefore, $f_{S}\in L^{\infty}(\mathbb{R}_{+};L^{p}(\Omega\times V))$. ∎
Now, we prove the uniqueness of the solution.
Theorem 3.
Consider the same conditions of Thm. 2, with $p=2$. Then the solution given by Thm. 2 is unique.
Proof.
We assume again that $C$ is an arbitrary, possibly distinct, constant. Let us define the norm
$\|\mathbf{f}-\mathbf{g}\|^{2}:=\int_{\Omega\times V}\left[(f_{S}-g_{S})^{2}+(f_{I}-g_{I})^{2}+(f_{R}-g_{R})^{2}\right]\mathrm{d}v\,\mathrm{d}x\ .$ (19)
From the linearity of $\mathcal{Q}^{(R)}_{j}$ it is not difficult to prove that
$\mathcal{Q}^{(R)}_{j}[\mathbf{f}]-\mathcal{Q}^{(R)}_{j}[\mathbf{g}]=\mathcal{Q}^{(R)}_{j}[\mathbf{f}-\mathbf{g}]\ .$
@@
$\int_{\Omega\times V}\left(\mathcal{Q}^{(R)}_{j}[\mathbf{f}]-\mathcal{Q}^{(R)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\mathcal{Q}^{(R)}_{j}[\mathbf{f}-\mathbf{g}](f_{j}-g_{j})\mathrm{d}v\,\mathrm{d}x$
$\leq 0\ .$ (20)
@@
$\left(\mathcal{Q}^{(T)}_{S}[\mathbf{f}]-\mathcal{Q}^{(T)}_{S}[\mathbf{g}]\right)\left(f_{S}-g_{S}\right)+\left(\mathcal{Q}^{(T)}_{I}[\mathbf{f}]-\mathcal{Q}^{(T)}_{I}[\mathbf{g}]\right)\left(f_{I}-g_{I}\right)$
$\qquad+\left(\mathcal{Q}^{(T)}_{R}[\mathbf{f}]-\mathcal{Q}^{(T)}_{R}[\mathbf{g}]\right)\left(f_{R}-g_{R}\right)$
@@
Using the inequality for $\mathcal{Q}^{(T)}_{j}$ and the classical inequality $ab\leq\frac{1}{2}(a^{2}+b^{2})$, we prove
$\sum_{j=S,I,R}\left(\mathcal{Q}^{(T)}_{j}[\mathbf{f}]-\mathcal{Q}^{(T)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)$
$\quad\leq C_{1}(f_{S}-g_{S})^{2}+C_{2}(f_{I}-g_{I})^{2}+C_{3}(f_{R}-g_{R})^{2}\ .$
Finally, we conclude that
$\int_{\Omega\times V}\sum_{j=S,I,R}\left(\mathcal{Q}^{(T)}_{j}[\mathbf{f}]-\mathcal{Q}^{(T)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)\mathrm{d}v\,\mathrm{d}x\leq C\|\mathbf{f}-\mathbf{g}\|^{2}\ .$ (21)
Finally, from Eqs. (20), (21), and the definition of $\mathfrak{Q}_{j}$, we conclude that
$\frac{1}{2}\frac{\mathrm{d}\ }{\mathrm{d}t}\|\mathbf{f}-\mathbf{g}\|^{2}$ $=\sum_{j}\int_{\Omega\times V}(f_{j}-g_{j})\left(\partial_{t}(f_{j}-g_{j})+v\cdot\nabla_{x}(f_{j}-g_{j})\right)\mathrm{d}v\,\mathrm{d}x$
$=\sum_{j}\int_{\Omega\times V}\mathfrak{Q}_{j}[\mathbf{f}-\mathbf{g}](f_{j}-g_{j})\mathrm{d}v\,\mathrm{d}x$
$\leq C\sum_{j}\int_{\Omega\times V}(f_{j}-g_{j})^{2}\mathrm{d}v\mathrm{d}x=C\|\mathbf{f}-\mathbf{g}\|^{2}\ ,$
and the uniqueness follows from an standard application of Gronwall’s inequality for initial conditions $\mathbf{f}^{\mathrm{ini}}=\mathbf{g}^{\mathrm{ini}}$. ∎
@@
As discussed at the beginning of this section, we did not consider the preferred kernel in the proof of Thms. 2 and 3. As will be discussed in the conclusions, we believe that the existence and uniqueness theorem still holds with parameters $\eta,\xi>0$.
@@
5 Examples and applications
The primary motivation for incorporating kinetic modelling into epidemiology is to elucidate how variations in population motility influence disease transmission [BBDP21]. This approach is especially pertinent when policymakers evaluate large-scale confinement strategies. Comprehensive reviews of the effects of widespread prophylactic measures on disease dynamics are available in [Law24, KDGM25].
A systematic analysis of the effects of confining I individuals would require considering class-dependent admissible velocity sets $V$, which extend beyond the current theoretical framework. Nevertheless, the effects of confining I individuals in one-dimensional cases will be formally addressed in Subsec. 5.4.
Before that, we will consider three other examples of kinetic models that are consistent with the SIRS epidemic model. In Subsec. 5.1, we will simulate the kinetic models numerically to illustrate how individuals disperse within the discretised mesoscopic model.
The second example, which will be developed in Subsec. 5.2, is pedagogical in nature and deals with a situation involving constant parameters, in order to clearly demonstrate the relationship between the kinetic model and the standard (space-homogeneous) SIRS dynamics. This is also important for understanding further expansions of our theory to include more general interaction kernels, which will be discussed in the conclusions.